OpenStudy (anonymous):

integral of sin(x^1/2)

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin \sqrt{x}\] to make it clear and \[u^2 = x\] is the substitution

OpenStudy (anonymous):

Do you know integral by parts?

OpenStudy (anonymous):

u'v+uv'?

OpenStudy (anonymous):

In this case, you should take substitution first, and then by parts.

OpenStudy (anonymous):

how did u conclude to that?

OpenStudy (anonymous):

Let t = \(\sqrt x\) --> dt=\(\dfrac{1}{2\sqrt x}dx\) you have \(2\sqrt x dt=dx\) or 2tdt = dx

OpenStudy (anonymous):

the integral becomes \( \int 2tsint dt\) now, apply integral by parts

OpenStudy (anonymous):

let u = 2t, dv = sint dt you can handle from here, right?

OpenStudy (anonymous):

yes thanx:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[2\sin \sqrt{x}-2\sqrt{x}\cos \sqrt{x}\] is what I got which is correct also according to the book.

OpenStudy (anonymous):

Now, how did you come to this conclusion on deciding that what should be dt and dx? alot of practice or intuition?

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

The letter I used, t, for the first substitute is just for not confused with the second part. If I use u = sqrt x, It will be unclear on the second part when we take integral by parts. That's it.

OpenStudy (anonymous):

okey, thanx for making it clear:)